OpenStudy (anonymous):

Solve: (x - 8)(x + 11) = 0. {8, -11} {-8, 11} {0, 11/8} no solution 9. Solve the equation (4x + 1)(3x - 2) = 0 {-4, 1.5} {-1/4, 2/3} {1/4, 2/3} {-4, 3}

OpenStudy (anonymous):

For the first one, equate x-8=0 or x+11=0 For the second, equate 4x+1=0 or 3x-2=0

OpenStudy (anonymous):

So, for the first when x-8=0 --> x=0+8 --> x=8 or solve the second one, and the answer should be the first one on the list. Try the second, and see if it works

OpenStudy (anonymous):

second one.. 4x+1= 0 or 3x-2=0 4x=-1 3x=2 x=-1/4 x=2/3